3. At time t-0 a particle is represented by the wave function A-if 0 < x<a ψ(x,0) = 0 otherwise where A, a, and b are constants. a) Normalize ψ(x,0). b) Draw (x,0). c) Where is the particle most likely to be found at t-0? d) What is the probability of finding the particle to the left of a? e) What is the expectation value of x?
Answer #1

Normaliz the function. 1=Lyydx dr - x -x Jab- ab-a - II A 1= (A *19 v*d+(64116=xtºck 1- (A) prd={A} [(6+2 -26)dt T0-03) (02.16) (x,0) X Consider A, a, and b are positive constants. The above diagram is the plot of function v (0,x). Since the slope o